# Statistical Inferences and Predictions for Areal Data and Spatial Data Fusion with Hausdorff–Gaussian Processes

### v2

@@ 4.1 Areal Data
We conducted a simulation study to assess the performance of the HGP relative to the DAGAR model, a state-of-the-art model designed explicitly for areal data (Datta et al., 2019), in modeling data generated from other models. We grounded our simulations in a real-world context by utilizing the 134 Intermediate Zones (IZ) north of the River Clyde within the Greater Glasgow in Scotland (Figure LABEL:fig:sm:areal_data), mirroring the application detailed in the upcoming Section 5.1. This region exhibits a range of Hausdorff distances between zones, from 0.7 to 37.7 kilometers (median: 9.4 km). Data were generated under a Poisson GLMM with a log link function, incorporating a single explanatory variable drawn from a standard normal distribution. The intercept and regression coefficients associated with this explanatory variable were fixed at 4.28 and 0.33, respectively, to mirror the parameters estimated in the real-world application (Section 5.1).
Two random effect structures were considered: the aggregated Gaussian process (AGP) and the Besag–York–Mollié (Besag et al., 1991, BYM), with model parameters corresponding to their fits to the real data. The AGP assumes a point-level GP aggregated at the areal unit level and is closer in structure to the HGP model. We chose AGP’s model parameters to reflect the HGP results from our real-world application (Section 5.1): a point-level marginal standard deviation ($\sigma$) of 0.19, a smoothness parameter ($\nu$) of 1, and a range parameter ($\rho$) of 2.21. On the other hand, the BYM model represents a spatially correlated random effect closer to the structure of the DAGAR model. The BYM model was conveniently parametrized by the marginal precision (Riebler et al., 2016). We set the marginal standard deviation ($\sigma$) to 0.24 and the spatial dependence parameter ($\zeta$) to 0.43, aligning with estimates from the real data analyses in Section 5.1. Section LABEL:sec:sm:sim of the appendix provides further details on the models used to generate data.
We generated 200 datasets from the Poisson GLMM under each of the two distinct random effect structures. For each simulated dataset, we fit two Bayesian Poisson GLMMs, the first one using an homoscedastic HGP while the second one using a DAGAR prior for the random effects. Both specified the fixed effects correctly with the single covariate. The priors for the parameters in the HGP model were set to be those presented in Section 3.2. The priors for the parameters in the DAGAR model were set to the same as those in(Datta et al., 2019). In particular, we set $\pi(\beta_{0})\propto 1$ and $\beta_{1}\sim N(0,1000^{2})$ for the regression parameters; a Gamma prior with shape 2 and rate 1 was placed on the conditional precision ($\phi$), and a uniform prior between 0 and 1 on the spatial dependence parameter ($\psi$). The GoF was evaluated using the criteria in Section 3.2, namely, LOOIC and WAIC, to compare which model fits the data better.
@@
Figure 2 shows the HGP consistently fitting the data better than the DAGAR model when data were generated from the AGP. The figure displays empirical cumulative distribution functions (ECDFs) of the differences ($\Delta_{G}$) in LOOIC and WAIC (left and right columns of the panel, respectively) between the models. While no formal thresholds exist for significant differences in the model comparison criteria we used (Vranckx et al., 2021), we consider differences exceeding 4 noteworthy. For AGP-generated data (solid line), our simulation study provides strong evidence that the HGP outperforms the DAGAR in GoF. Specifically, $\Delta_{\textrm{LOOIC}}$ was below 10-10 for 80% of simulated datasets, strongly favoring the HGP. In contrast, evidence favoring DAGAR while exceeding ($\Delta_{\textrm{LOOIC}}>4$) was observed in only 1% of the generated data. Interestingly, the HGP remained competitive even when we generated data from the BYM model (a structure closer to the DAGAR). While there was some evidence favoring the DAGAR, $\Delta_{\textrm{LOOIC}}$ was below 4 for more than 50% of the simulations under this scenario. These results highlight the HGP’s robustness as a disease mapping model, demonstrating superior performance when the data aligns with its structure and remains a viable option even when a specialized competitor better represents the underlying process.
In the appendix (Table LABEL:tab:sm:areal), we also evaluated the HGP’s ability to recover true model parameters under correct specification. We assessed estimation accuracy using Mean Absolute Percentage Error (MAPE) and the frequentist coverage percentage (CP) of 95% credible intervals. All model parameters achieved CPs close to the nominal level, with the lowest being 92.5% for $\rho$, which consistently proved the most challenging to estimate. The second most difficult parameter, $\sigma$, exhibited a maximum MAPE of only 7%.
@@ 4.2 Data Fusion
To assess the predictive performance of the HGP against standard data fusion models, we conducted a simulation study mirroring usual data fusion scenarios. Within a square region with an area of $51^{2}$, we generated a map comprising a $17\times 17$ grid-boxes of areal units (each with an area of $3^{2}$) and 35 point-referenced data points. Additionally, we created five extra spatial units for each of the following three types of spatial units used exclusively to evaluate predictions: larger polygons (grid boxes with area of $3^{2}$), smaller polygons (grid boxes with area of $1.5^{2}$), and point-referenced, as displayed in Figure 3. We carefully chose the study region and grid box dimensions so each polygon’s area exceeded one, aligning with typical data fusion problems. This design emulates scenarios where areal units, often derived from satellite or computer models, are more abundant than point-referenced data. Hausdorff distances between sample units in this map ranged from 0.4 to 69.8. We simulate the datasets from a Normal GLMM with an identity link function, no covariates, and a spatial random effect. The overall mean of this GLMM was fixed at 6.28 to mirror the application from Section 5.2. In addition, this model introduces the small scale variance (denoted $\tau^{2}$).
[…]
@@
where $\{Z(\mathbf{s})\,:\,\mathbf{s}\in D\}$ is a zero-mean, second-order stationary GP with an isotropic Matérn covariance function with smoothness parameter $\nu=1$. SeeMoraga et al. (2017) for a detailed definition of this function. We simulated AGP data with parameters $\sigma^{2}=2.71^{2}$, $\rho=2.29$, and $\tau^{2}=1\times{10}^{-16}$, again mirroring the real data analysis in Section 5.2. In aggregated models, $\tau$ approaches zero as only the average of homoscedastic independent residuals is observed, which converges in probability to zero.
@@
We then generated 200 datasets from the Normal GLMM under each of the two different spatial random effects’ structures detailed next. For each simulated dataset, we fit two Bayesian Normal GLMMs, one using the aforementioned heteroscedastic HGP prior for the random effect and the other using the AGP prior. When fitting the HGP model, we used the priors highlighted in Section 3.2. In addition, we put a penalized complexity prior (Simpson et al., 2017) on $\tau^{-2}$, such that $\mathbbm{P}(\tau>a)=.05$, where $a$ is half the empirical SD observed in the data. Inferences and predictions were based in 2000 MCMC samples with a 1000 warm-up phase per chain with four parallel chains. The initialization of the algorithm is described in Section 3.2.
@@
Inferences for the model induced by the spatial random effect from Equation (9) were performed using the integrated nested Laplace approximation and the stochastic partial differential equation (Lindgren et al., 2011, SPDE) method. The joint distribution induced by this AGP involve approximating stochastic integrals. Practitioners must define a grid with a fixed number of points within each areal unit to approximate these integrals (Wang and Furrer, 2021). The SPDE uses a mesh to discretize the space to achieve such an approximation efficiently. The number of grid points per polygon will depend on the resolution of the mesh. We used two different meshes to show the inferences and, consequently, the predictions may change substantially depending on the accuracy of the approximation for the stochastic integrals in Equation (9). In the appendix , Figure LABEL:fig:sm:meshes-sim visualizes these meshes: a sparse mesh (left column), and a fine-resolution mesh (second column), with approximately 2 and 25 points per polygon, respectively. Henceforth, we will differentiate the AGP models by mesh density, denoting the sparse-mesh model as AGP1 and the fine-mesh model as AGP2. AGP model priors are detailed in Section LABEL:sec:sm:add-sim-fused of the appendix .
@@
We assess the quality of model predictions across different scenarios using four metrics. Firstly, point and interval predictions are derived from the posterior predictive distributions (Equation (8)), using their median and the 2.5% and 97.5% percentiles, respectively. The root mean square error of prediction (RMSP) is used to evaluate point estimates. The remaining three metrics assess interval predictions: the frequentist coverage percentage of prediction intervals, denoted CPP, their width, and the interval score (Gneiting and Raftery, 2007, IS). The latter is defined as
@@
Table 1: Out-of-sample prediction for the three models under different data generating models. Spatial unit indicates the geometry of the sample units for which predictions were assessed; for polygons, area is additionally specified within parenthesis. Fit specifies the model fitted to the data. RMSP is the root mean squared error of prediction, CPP is the frequentist coverage percentage of the 95% prediction interval, Width is the prediction interval width, and IS is the interval score. Double headers denote the true data-generating model.
@@
Spatial unit Fit RMSP CPP Width IS RMSP CPP Width IS
@@
Table 1 further reveals demonstrates a consistent decline in prediction accuracy for all models as the area of the spatial units decreased. For AGP-generated data, the AGP2 model’s RMSP for point-referenced data was approximately four times higher than for the polygons with bigger area (i.e., $3^{2}$). Similarly, the HGP model’s RMSP increased by a factor of 3.15 for the same comparison (point-referenced vs largest polygon). For both the HGP and AGP2, CPP remained near the nominal level for all the spatial units. However, HGP’s CPP dropped below nominal level for smaller polygons (area of $1.5^{2}$). This is likely due to variance underestimation: while the AGP-generated data’s variance is inversely proportional to spatial units, the fitted HGP model only distinguish between points and polygons, not their area. Regarding the IS, both our model and the AGP2 model showed IS values roughly three times higher for point-referenced data compared to the largest polygons. This aligns with the simulation parameters. The AGP1 model’s IS performance deteriorated substantially when decreasing the area of the spatial units.
Under the HGP data-generating process, the HGP model consistently outperformed competitors on all assessment metrics except PI width. Although point predictions were broadly similar across models, the HGP provided remarkably superior interval predictions for all spatial unit types, mantaining CPP near the nominal level (see Table 1. In contrast, the AGP1 model exhibited a CPP of only 22.8% for point-referenced data. The poor coverage for both AGP models with point-referenced data is further reflected in their IS, which were 5.9 (AGP1) and 1.6 (AGP2) times higher than the HGP. On the other hand, the AGP2 model remained competitive with the HGP for polygons interval predictions, while the sparse AGP performed poorly.
Table LABEL:tab:sm:fusion in the appendix presents the results for parameter estimation under the correct specification of the the HGP model. We used the same metrics as in Section 4.1 to assess the parameters estimation. The MAPE for $\rho$ was around $10.9\%$ while for $\sigma_{a}$ it was $9.21\%$. Similarly to what is observed with geostatistical models, the small scale standard deviation parameter, namely $\tau$, was the most challenging parameter to be estimated. Unlike the areal simulation, the coverage of some of the variance parameters was consistently below the nominal level of $95\%$. This results are expected, as the HGP inherits the weak identifiability of these parameters from geostatistical models under a normal likelihood (Zhang, 2004). Predictions remain trustworthy despite this issue.
@@ 5.1 Areal Data: Respiratory Disease Hospitalization
We first analyze a dataset on the count of hospital admissions due to respiratory disease in 2010 for the 134 intermediate zones (IZ) to the north of the river Clyde in the Great Glasgow and Clyde health board in Scotland. This dataset is publicly available in theCARBayesdata package (Lee, 2022). For each IZ, the expected number of admissions based on the national age- and sex-standardized rates, alongside the percentage of people classified as income deprived in each IZ are also available. The standardized morbidity ratio (SMR), the ratio between the number of observed and expected cases, is often used to control for confounding factors when estimating risks (Greenland, 1984). The goals of the analyses are: (1) to provide smooth estimates of the SMR, adjusted for the percentage of income deprivation, and (2) to contrast the outcomes from the proposed methodology with other established methods.
@@
In this application, $Y(\mathbf{s}_{i})$ is the number of admissions by respiratory disease registered at the $i$-th IZ, denoted $\mathbf{s}_{i}$. In the $i$-th IZ, define $x_{i}$ as the percentage of people who are defined to be income deprived and $E_{i}$ as the expected number of admissions. We analyze the data using the GLMM model from Section 3.1, with a Poisson likelihood, a $\log$ link function, an intercept, a covariate $x_{i}$, and an offset $E_{i}$. We employed the homoscedastic HGP (with a PEC function with smoothness $\nu=0.7$) as the prior for the random effects; the choice of $\nu=0.7$ is justified in Table LABEL:tab:sm:gof_scot of the appendix. The priors for the parameters in the HGP GLMM were set to be those presented in Section 3.2. Inference was performed using four parallel chains with 5000 MCMC samples, thinned by 10, after discarding the initial 1000 iterations as the warm-up period. The initialization of the algorithm is described in Section 3.2. The highest split-$\hat{R}$ was smaller than $1.01$, providing evidence of convergence.
@@
Two competing models, BYM and DAGAR, were considered as the priors for the spatial random effects $\mathbf{z}$ for comparison, as outlined in Sections 4.1 and LABEL:sec:sm:sim, respectively. In the BYM model, parametrized by the marginal precision (Riebler et al., 2016), we followedMorris et al. (2019) and set standard normal priors for the intercept ($\beta_{0}$), regression coefficient ($\beta_{1}$), and the $\mathrm{logit}$ of the mixing parameter $\zeta$. Similarly, a standard normal distribution truncated on positive real numbers was used for $\sigma$. The priors for the DAGAR model are the ones presented in Section 4.1. The initialization of the MCMC algorithm followed the same principle used for the HGP model. In both cases, we used four parallel chains with a warm-up period of 4000 iterations and based the inferences on 20000 MCMC samples, thinned by 20. For these models, the parameters split-$\hat{R}$ provided evidence of convergence. In addition, trace and density plots for posterior samples of all fitted model parameters are provided in Section LABEL:sec:sm:areal_convergence of the appendix.
@@
We present the posterior medians and 95% CIs for the parameters from the HGP, DAGAR, and BYM models in Table 2. When considering the coefficient linked to the covariate representing income deprivation, all three models yield similar posterior estimates. The positive posterior estimates the respective parameter ($\beta_{1}$) indicate a positive association between proportion of income deprived people and hospitalization rates for respiratory diseases. Among these models, DAGAR exhibits the broadest credible intervals for the intercept, whereas the marginal posterior distribution for this parameter in the BYM model is highly concentrated around 0.22-0.22. Examining the spatial aspect, the confidence interval for the practical range parameter $\rho$ in the HGP model implies that the distance at which the spatial correlation between random effects from two distinct spatial units decreases to $0.10$ is at most $6.95$ kilometers. Conversely, the 95% CI for the spatial dependence parameter $\psi$ from the DAGAR model spans from $0.07$ to $0.83$. In other words, this credible interval suggests that the clustering effect ranges from almost negligible to very strong. A similar pattern is observed for the $\zeta$ in the BYM model.
@@
To gain a better understanding of this value in the context of the data, Figure 4 displays a density histogram of all pairwise Hausdorff distances observed in the dataset. The plot also features the prior distribution (in red) and estimated posterior density (in blue) associated with the practical range. Notably, the majority of the posterior distribution mass for $\rho$ falls below the concentration of most pairwise distances, indicating a negligible clustering effect in hospital admissions. An additional analysis of the spatial correlation is presented in Section LABEL:sec:sm:areal_comp in the appendix .
In Table 2, we also present two model comparison criteria described in Section 4 for the models fitted to the data. These results suggest the HGP stands out as the best fit based on the LOOIC, rendering the HGP prior a competitive fit for these data. Despite the modest improvement in model fitting capacity, the HGP possesses an advantage in terms of intuitive interpretability of its spatial dependence, achieved in the last two paragraphs through the practical range parameter $\rho$. Among the specialized models for areal data, the BYM was the least adequate in this application. However, none of the models delivered a remarkably better fit to this dataset. This is further reinforced by looking at the maps of observed and smoothed SMR estimates presented in Figure LABEL:fig:sm:smr from the appendix.

### v3

@@ 4.1 Areal Data
We conducted a simulation study to assess the performance of the HGP relative to the DAGAR model, a state-of-the-art model designed explicitly for areal data (Datta et al.,, 2019), in modeling data generated from other models. We grounded our simulations in a real-world context by utilizing the 134 Intermediate Zones (IZ) north of the River Clyde within the Greater Glasgow in Scotland (Figure S.1), mirroring the application detailed in the upcoming Section 5.1. This region exhibits a range of Hausdorff distances between zones, from 0.7 to 37.7 kilometers (median: 9.4 km). Data were generated under a Poisson GLMM with a log link function, incorporating a single explanatory variable drawn from a standard normal distribution. The intercept and regression coefficients associated with this explanatory variable were fixed at 4.28 and 0.33, respectively, to mirror the parameters estimated in the real-world application (Section 5.1).
Two random effect structures were considered: the aggregated Gaussian process (AGP) and the Besag–York–Mollié (Besag et al.,, 1991, BYM), with model parameters corresponding to their fits to the real data. The AGP assumes a point-level GP aggregated at the areal unit level and is closer in structure to the HGP model. We chose AGP’s model parameters to reflect the HGP results from our real-world application (Section 5.1): a point-level marginal standard deviation ($\sigma$) of 0.19, a smoothness parameter ($\nu$) of 1, and a range parameter ($\rho$) of 2.21. The BYM model allows for spatially correlated random effects with a structure closer to that of the DAGAR model. We set its marginal standard deviation (Riebler et al.,, 2016), denoted $\sigma$, to 0.24 and the spatial dependence parameter ($\xi$) to 0.43, aligning with estimates from the real data analyses in Section 5.1. Section S.2.1 of the Appendix Material provides further details on the models used to generate data.
We generated 200 datasets from the Poisson GLMM under each of the two distinct random effect structures. For each simulated dataset, we fit two Bayesian Poisson GLMMs, the first one using an homoscedastic HGP while the second one using a DAGAR prior for the random effects. Both specified the fixed effects correctly with the single covariate. The priors for the parameters in the HGP model were set to be those presented in Section 3.2. The priors for the parameters in the DAGAR model were set to the same as those in(Datta et al.,, 2019). In particular, we set $\pi(\beta_{0})\propto 1$ and $\beta_{1}\sim N(0,1000^{2})$ for the regression parameters; a Gamma prior with shape 2 and rate 1 was placed on the conditional precision ($\phi$), and a uniform prior between 0 and 1 on the spatial dependence parameter ($\psi$). The GoF was evaluated using the criteria in Section 3.2, namely, LOOIC and WAIC, to compare which model fits the data better.
@@
Figure 3 shows the HGP consistently fitting the data better than the DAGAR model when data were generated from the AGP. The figure displays empirical cumulative distribution functions (ECDFs) of the differences ($\Delta_{G}$) in LOOIC and WAIC (left and right columns of the panel, respectively) between the models. While no formal thresholds exist for significant differences in the model comparison criteria we used, we consider differences exceeding 5 to be noteworthy, following a similar practice used with the DIC (Vranckx et al.,, 2021). For AGP-generated data (solid line), our simulation study provides strong evidence that the HGP outperforms the DAGAR in GoF. Specifically, $\Delta_{\textrm{LOOIC}}$ was below 10-10 for 80% of simulated datasets, strongly favoring the HGP. In contrast, evidence favoring DAGAR ($\Delta_{\textrm{LOOIC}}>4$) was observed in only 1% of the generated data. Interestingly, the HGP remained competitive even when we generated data from the BYM model (a structure closer to the DAGAR). While there was some evidence favoring the DAGAR, $\Delta_{\textrm{LOOIC}}$ was below 4 for more than 50% of the simulations under this scenario. These results highlight the HGP’s robustness as a disease mapping model, demonstrating superior performance when the data aligns with its structure and remains a viable option even when a specialized competitor better represents the underlying process.
In the Appendix Material (Table S.1), we also evaluated the HGP’s ability to recover true model parameters under correct specification. We assessed estimation accuracy using Mean Absolute Percentage Error (MAPE) and the frequentist coverage percentage (CP) of 95% credible intervals. All model parameters achieved CPs close to the nominal level, with the lowest being 92.5% for $\rho$, which consistently proved the most challenging to estimate. The marginal standard deviation ($\sigma$) exhibited a maximum MAPE of only 7%.
@@ 4.2 Data Fusion
To assess the predictive performance of the HGP against standard data fusion models, we conducted a simulation study mirroring usual data fusion scenarios. Within a square region with an area of $51^{2}$, we generated a map comprising a $17\times 17$ grid-boxes of areal units (each with an area of $3^{2}$) and 35 point-referenced data points. Additionally, we created five extra spatial units for each of the following three types exclusively to evaluate predictions: larger polygons (grid boxes with area of $3^{2}$), smaller polygons (grid boxes with area of $1.5^{2}$), and point-referenced, as displayed in Figure 4. We carefully chose the study region and grid box dimensions so each polygon’s area exceeded one, aligning with typical data fusion problems. This design emulates scenarios where areal units, often derived from satellite or computer models, are more abundant than point-referenced data. Hausdorff distances between sample units in this map ranged from $0.4$ to $69.8$. We simulate the datasets from a Normal GLMM with an identity link function, no covariates, and a spatial random effect. The overall mean of this GLMM was fixed at $6.28$ to mirror the application from Section 5.2. In addition, this model introduces the small scale variance (denoted $\tau^{2}$).
This section’s spatial GLMMs employ two distinct spatial random effect structures: a heteroscedastic HGP and an AGP. The HGP’s SD function $v(\mathbf{s})$ is defined as $\exp\{\alpha_{0}+\alpha_{1}\mathbbm{1}(\lvert\mathbf{s}\rvert>0)\}$, allowing for distinct variances between point-referenced and areal data. We selected HGP parameters ($\tau^{2}=0.17^{2}$, $\sigma=4.4416$, $\sigma_{a}=1.3114$, $\rho=28.45$, and $\nu=0.8$) to closely resemble the data analysis presented in Section 5.2. Note that the spatial variance for point and areal units ($\sigma$ and $\sigma_{a}$, respectively) is a one-to-one transformation of $(\alpha_{0},\alpha_{1})$$^{\top}$. We set the dependence parameter $\rho$ to 48.9. This value differs from our estimate in the real data due to a necessary rescaling based on the map used in this simulation study. For further details, refer to Section S.3 in the Appendix Material.
@@
where $\{Z(\mathbf{s})\,:\,\mathbf{s}\in D\}$ is a zero-mean, second-order stationary GP with an isotropic Matérn covariance function with smoothness parameter $\nu=1$. SeeMoraga et al., (2017) for a detailed definition of this function. We simulated AGP data with parameters $\sigma^{2}=2.71^{2}$, $\rho=2.29$, and $\tau^{2}=1\times{10}^{-16}$, again mirroring the real data analysis in Section 5.2. In aggregated models, $\tau$ approaches zero as only the average of homoscedastic independent residuals is observed, which converges in probability to zero.
@@
We then generated 200 datasets from the Normal GLMM under each of the two different spatial random effects’ structures detailed next. For each simulated dataset, we fit two Bayesian Normal GLMMs, one using the aforementioned heteroscedastic HGP prior for the random effect and the other using the AGP prior. When fitting the HGP model, we used the priors highlighted in Section 3.2. In addition, we put a penalized complexity prior (Simpson et al.,, 2017) on $\tau^{-2}$, such that $\mathbbm{P}(\tau>a)=.05$, where $a$ is half the empirical SD observed in the data. Inferences and predictions were based in 2000 MCMC samples with a 1000 warm-up phase per chain with four parallel chains. The initialization of the algorithm is described in Section 3.2.
@@
Inferences for the model induced by the spatial random effect from Equation (9) were performed using the integrated nested Laplace approximation and the stochastic partial differential equation (Lindgren et al.,, 2011, SPDE) method. The joint distribution induced by this AGP involve approximating stochastic integrals. Practitioners must define a grid with a fixed number of points within each areal unit to approximate these integrals (Wang and Furrer,, 2021). The SPDE uses a mesh to discretize the space to achieve such an approximation efficiently. The number of grid boxes per polygon will depend on the resolution of the mesh. We used two different meshes to show the inferences and, consequently, the predictions may change substantially depending on the accuracy of the approximation for the stochastic integrals in Equation (9). In the Appendix Material, Figure S.2 visualizes these meshes: a sparse mesh (left column), and a fine-resolution mesh (second column), with approximately 2 and 25 points per polygon, respectively. Henceforth, we will differentiate the AGP models by mesh density, denoting the sparse-mesh model as AGP1 and the fine-mesh model as AGP2. Their priors are detailed in Section S.4 of the Appendix Material.
@@
We assess the quality of model predictions across different scenarios using four metrics. Firstly, point and interval predictions are derived from the posterior predictive distributions (Equation (8)), using their median and the 2.5% and 97.5% percentiles, respectively. The root mean square error of prediction (RMSP) is used to evaluate point estimates. The remaining three metrics assess interval predictions: the frequentist coverage percentage of prediction intervals, denoted CPP, their width, and the interval score (Gneiting and Raftery,, 2007, IS). The latter is defined as
@@
Table 1: Out-of-sample prediction for the three models under different data generating models. “Sp. unit” indicates the geometry of the sample units for which predictions were assessed; for polygons (denoted Poly), area is additionally specified within parenthesis. Fit specifies the model fitted to the data. RMSP is the root mean squared error of prediction, CPP is the frequentist coverage percentage of the 95% prediction interval, Width is the prediction interval width, and IS is the interval score. Double headers denote the true data-generating model.
@@
Sp. unit Fit RMSP CPP Width IS RMSP CPP Width IS
@@
Table 1 further reveals a consistent decline in prediction accuracy for all models as the area of the spatial units decreased. For AGP-generated data, the AGP2 model’s RMSP for point-referenced data was approximately four times higher than for the polygons with bigger area (i.e., $3^{2}$). Similarly, the HGP model’s RMSP increased by a factor of 3.15 for the same comparison (point-referenced vs largest polygon). For both the HGP and AGP2, CPP remained near the nominal level for all the spatial units. However, HGP’s CPP dropped below nominal level for smaller polygons (area of $1.5^{2}$). This is likely due to variance underestimation: while the AGP-generated data’s variance is inversely proportional to spatial units, the fitted HGP model only distinguish between points and polygons, not their area. Regarding the IS, both our model and the AGP2 model showed IS values roughly three times higher for point-referenced data compared to the largest polygons. This aligns with the simulation parameters. The AGP1 model’s IS performance deteriorated substantially when decreasing the area of the spatial units.
Under the HGP data-generating process, the HGP model consistently outperformed competitors on all assessment metrics except PI width. Although point predictions were broadly similar across models, the HGP provided remarkably superior interval predictions for all spatial unit types, mantaining CPP near the nominal level (see Table 1). In contrast, the AGP1 model exhibited a CPP of only 22.8% for point-referenced data. The poor coverage for both AGP models with point-referenced data is further reflected in their IS, which were 5.9 (AGP1) and 1.6 (AGP2) times higher than the HGP. On the other hand, the AGP2 model remained competitive with the HGP for polygons interval predictions, while the sparse AGP performed poorly.
Table S.3 in the Appendix Material presents the results for parameter estimation under the correct specification of the the HGP model. We used the same metrics as in Section 4.1 to assess the parameters estimation. The MAPE for $\rho$ was around $10.9\%$ while for $\sigma_{a}$ it was $9.21\%$. Similarly to what is observed with geostatistical models, the small scale standard deviation parameter, namely $\tau$, was the most challenging parameter to be estimated. Unlike the areal simulation, the coverage of some of the variance parameters was consistently below the nominal level of $95\%$. This results are expected, as the HGP inherits the weak identifiability of these parameters from geostatistical models under a normal likelihood (Zhang,, 2004). Predictions remain trustworthy despite this issue.
@@ 5.1 Areal Data: Respiratory Disease Hospitalization
We first analyze a dataset on the count of hospital admissions due to respiratory disease in 2010 for the 134 intermediate zones (IZ) to the north of the river Clyde in the Great Glasgow and Clyde health board in Scotland. This dataset is publicly available in theCARBayesdata package (Lee,, 2022). For each IZ, the expected number of admissions based on the national age- and sex-standardized rates, alongside the percentage of people classified as income deprived in each IZ are also available. The standardized morbidity ratio (SMR), the ratio between the number of observed and expected cases, is often used to control for confounding factors when estimating risks (Greenland,, 1984). The goals of the analyses are: (1) to provide smooth estimates of the SMR, adjusted for the percentage of income deprivation, and (2) to contrast the outcomes from the proposed methodology with other established methods.
@@
In this application, $Y(\mathbf{s}_{i})$ is the number of admissions by respiratory disease registered at the $i$-th IZ, denoted $\mathbf{s}_{i}$. In the $i$-th IZ, define $x_{i}$ as the percentage of people who are defined to be income deprived and $E_{i}$ as the expected number of admissions. We analyze the data using the GLMM model from Section 3.1, with a Poisson likelihood, a $\log$ link function, an intercept, a covariate $x_{i}$, and an offset $E_{i}$. We employed the homoscedastic HGP (with a PEC function with smoothness $\nu=0.7$) as the prior for the random effects; the choice of $\nu=0.7$ is justified in Table S.4 of the Appendix Material. The priors for the parameters in the HGP GLMM were set to be those presented in Section 3.2. Inference was performed using four parallel chains with 5000 MCMC samples, thinned by 10, after discarding the initial 1000 iterations as the warm-up period. The initialization of the algorithm is described in Section 3.2. The highest split-$\hat{R}$ was smaller than $1.01$, providing evidence of convergence.
@@
Two competing models, BYM and DAGAR, were considered as the priors for the spatial random effects $\mathbf{z}$ for comparison, as outlined in Sections 4.1 and S.2.1, respectively. In the BYM model, parametrized by the marginal precision (Riebler et al.,, 2016), we followed Morris et al., (2019) and set standard normal priors for the intercept ($\beta_{0}$), regression coefficient ($\beta_{1}$), and the $\mathrm{logit}$ of the mixing parameter $\xi$. Similarly, a standard normal distribution truncated on positive real numbers was used for $\sigma$. The priors for the DAGAR model are the ones presented in Section 4.1. The initialization of the MCMC algorithm followed the same principle used for the HGP model. In both cases, we used four parallel chains with a warm-up period of 4000 iterations and based the inferences on 20000 MCMC samples, thinned by 20. For these models, the parameters split-$\hat{R}$ provided evidence of convergence. In addition, trace and density plots for posterior samples of all fitted model parameters are provided in Section S.5.1 of the Appendix Material.
@@
We present the posterior medians and 95% CIs for the parameters from the HGP, DAGAR, and BYM models in Table 2. When considering the coefficient linked to the covariate representing income deprivation, all three models yield similar posterior estimates. The positive posterior estimates the respective parameter ($\beta_{1}$) indicate a positive association between proportion of income deprived people and hospitalization rates for respiratory diseases. Among these models, DAGAR exhibits the broadest credible intervals for the intercept, whereas the marginal posterior distribution for this parameter in the BYM model is highly concentrated around 0.22-0.22. Examining the spatial aspect, the confidence interval for the practical range parameter $\rho$ in the HGP model implies that the distance at which the spatial correlation between random effects from two distinct spatial units decreases to $0.10$ is at most $6.95$ kilometers. Conversely, the 95% CI for the spatial dependence parameter $\psi$ from the DAGAR model spans from $0.07$ to $0.83$. In other words, this credible interval suggests that the clustering effect ranges from almost negligible to very strong. A similar pattern is observed for the parameter $\xi$ in the BYM model.
@@
To gain a better understanding of this value in the context of the data, Figure 5 displays a density histogram of all pairwise Hausdorff distances observed in the dataset. The plot also features the prior distribution (in red) and estimated posterior density (in blue) associated with the practical range. Notably, the majority of the posterior distribution mass for $\rho$ falls below the concentration of most pairwise distances, indicating a negligible clustering effect in hospital admissions. An additional analysis of the spatial correlation is presented in Section S.5 in the Appendix Material.
In Table 2, we also present two model comparison criteria described in Section 4 for the models fitted to the data. These results suggest the HGP stands out as the best fit based on the LOOIC, rendering the HGP prior a competitive fit for these data. Despite the modest improvement in model fitting capacity, the HGP possesses an advantage in terms of intuitive interpretability of its spatial dependence, achieved in the last two paragraphs through the practical range parameter $\rho$. Among the specialized models for areal data, the BYM was the least adequate in this application. However, none of the models delivered a remarkably better fit to this dataset. This is further reinforced by looking at the maps of observed and smoothed SMR estimates presented in Figure S.3 from the Appendix Material.